OpenStudy (calculusxy):

Melissa drives from Central City to Marbury, a distance of 154 miles. If she leaves Central City at 8 A.M. and arrives in Marbury at 10:45 A.M., what is her average rate of speed?

OpenStudy (calculusxy):

@Mertsj

OpenStudy (texaschic101):

154 divided by 2 hrs and 45 minutes ?

OpenStudy (mertsj):

\[\frac{154}{2.75}=\frac{56mi}{1hr}\]

OpenStudy (calculusxy):

I tried that but it doesn't work @texaschic101

OpenStudy (texaschic101):

oh I see...I was wondering what to do with the 2 hrs and 45 minutes....

OpenStudy (mertsj):

2hr 45 min = 2 45/60 hr = 2 3/4 hr = 2.75 hrs

OpenStudy (calculusxy):

Thank you @texaschic101 and @Mertsj

OpenStudy (texaschic101):

I understand now...thx Mertsj

OpenStudy (mertsj):

yw